Joe drives 20 miles in a half hour. How long will it take him to drive 50 mi?
Found 5 solutions by checkley71, amalm06, Alan3354, josgarithmetic, TeachMath:
Answer by checkley71(8403)

1/2 HOUR = 20 MILES THEREFORE HE IS DRIVING @ 40 MPH.
SO IF HE DRIVES 50 MILES @ 40 MPH IT WILL TAKE HIM 50/40 OR 1.25 HOURS.

Answer by amalm06(224)

(30 min/20 mi)(50 mi)=75 min (Answer)

Answer by Alan3354(69443)

Will he speed up? Slow down?
If you mean at the same rate, you should say that.
---
Not sufficient info.

Answer by josgarithmetic(39620)
Answer by TeachMath(96)

20/50 = 0.5/T
20T = 50(.5)
T, or time = 1.25 hours